# Stabilizer Entanglement Distillation and Efficient Fault-Tolerant Encoder

### v1

@@ IV Decoding Stabilizer Code by Single-Qubit Measurements
We propose a decoder that transforms logical Bell states into physical Bell states using single-qubit measurements and conditional Pauli operators. Since stabilizer and single-qubit measurements can be performed in parallel, this significantly reduces the circuit depth for implementing stabilizer entanglement distillation protocols.
@@
To develop the decoder, we represent the stabilizer code using a binary check matrix $\left(\bm{X}\middle|\bm{Z}\right)$ [90], where $\bm{X}$ and $\bm{Z}$ are submatrices of dimensions $\left(n-k\right)\times n$. The rows represent generators $g_{1}$ through $g_{n-k}$, and the columns of each submatrix correspond to the $n$ physical qubits. In the $i$th row, represented by vector $\left(\bm{x}_{i}\middle|\bm{z}_{i}\right)$, $x_{ij}=1$ indicates a Pauli-$X$ on the $j$th qubit, $z_{ij}=1$ indicates a Pauli-$Z$, and $1$ on both indicates a Pauli-$Y$. Two generators $g_{i}$ and $g_{j}$ commute if and only if their symplectic inner product is zero, given by
@@
which follows the same dimensions. The $i$th row, $L_{Z}^{i}$, represents the $Z$ operator for the $i$th logical qubit. These logical operators commute with the stabilizer generators. $X_{L}^{i}$ and $Z_{L}^{i}$ anti-commute, while $X_{L}^{i}$ and $Z_{L}^{j}$ commute for $i\neq j$.
@@
    The logical operator $X_{L}^{i}$ is independent of $X_{j}$ for $j\leq r$ and $n-k<j<n-k+i$, and independent of $Z_{j}$ for $r<j\leq n$. The logical operators $Z_{L}^{i}$ is independent of all $X$s and independent of $Z_{j}$ for $r<j<n-k+i$.
@@
$\begin{matrix}g_{1}\otimes g_{1}\;,&Z_{1}\otimes Z_{1}\\ \vdots&\vdots\\ g_{r}\otimes g_{r}\;,&Z_{r}\otimes Z_{r}\\ X_{r+1}\otimes X_{r+1}\;,&g_{r+1}\otimes g_{r+1}\\ \vdots&\vdots\\ X_{n-k}\otimes X_{n-k}\;,&g_{n-k}\otimes g_{n-k}\\ X_{L}^{1}\otimes X_{L}^{1}\;,&Z_{L}^{1}\otimes Z_{L}^{1}\\ \vdots&\vdots\\ X_{L}^{k}\otimes X_{L}^{k}\;,&Z_{L}^{k}\otimes Z_{L}^{k}\\ \end{matrix}\;.$ (7)
@@
$X_{L}^{i}=\left(-1\right)^{\alpha_{x}^{i}}X_{n-k+i}\;,$ (10)
where $\alpha_{x}^{i}=\sum_{j:C_{ij}^{T}=1}a_{j}+\sum_{j:E_{ij}^{T}=1}a_{r+j}$. Similarly, the logical operator $Z_{L}^{i}$ is given by
$Z_{L}^{i}=\left(-1\right)^{\alpha_{z}^{i}}Z_{n-k+i}\;,$ (11)
where $\alpha_{z}^{i}=\sum_{j:\left(A_{2}^{T}\right)_{ij}=1}a_{j}$ and $\left(A_{2}^{T}\right)_{ij}$ is the element of submatrix $\bm{A}_{2}^{T}$ from Equation 5. The single-qubit measurements anti-commute with the code generators, updating the stabilizers in Equation 8 to
$\begin{matrix}\left(-1\right)^{a_{1}}Z_{1}\otimes I\;,&\left(-1\right)^{b_{1}}I\otimes Z_{r}\\ \vdots&\vdots\\ \left(-1\right)^{a_{r}}Z_{r}\otimes I\;,&\left(-1\right)^{b_{r}}I\otimes Z_{r}\\ \left(-1\right)^{a_{r+1}}X_{r+1}\otimes I\;,&\left(-1\right)^{b_{r+1}}I\otimes X_{r+1}\\ \vdots&\vdots\\ \left(-1\right)^{a_{n-k}}X_{n-k}\otimes I\;,&\left(-1\right)^{b_{n-k}}I\otimes X_{n-k}\\ \left(-1\right)^{\alpha_{x}^{1}+\beta_{x}^{1}}X_{n-k+1}\otimes X_{n-k+1}\;,&\left(-1\right)^{\alpha_{z}^{1}+\beta_{z}^{1}}Z_{n-k+1}\otimes Z_{n-k+1}\\ \vdots&\vdots\\ \left(-1\right)^{\alpha_{x}^{k}+\beta_{x}^{k}}X_{n}\otimes X_{n}\;,&\left(-1\right)^{\alpha_{z}^{k}+\beta_{z}^{k}}Z_{n}\otimes Z_{n}\\ \end{matrix}\;.$
Here, $\bm{b}=b_{1}b_{2}\ldots b_{n-k}$ denotes Bob’s decoding parities from single-qubit measurements, which are independent of Alice’s. $\beta_{x}^{i}$ and $\beta_{z}^{i}$ are computed in the same way as $\alpha_{x}^{i}$ and $\alpha_{z}^{i}$, respectively. After the measurements, Alice and Bob discard their first $n-k$ qubits and retain the last $k$ qubits, forming the decoded physical Bell states. They adjust the phases $\alpha_{x}^{i}$, $\alpha_{z}^{i}$, $\beta_{x}^{i}$, and $\beta_{z}^{i}$ to 00 by applying Pauli operators to the corresponding qubits, yielding $k$ standard Bell states.
Integrating stabilizer measurements, error correction, and our decoder results in the measurement-based stabilizer entanglement distillation protocol (see Appendix B for more details). The protocol is outlined as follows:
@@
This results in $k$ distilled Bell states shared between Alice and Bob. Note that two-way entanglement distillation, where the stabilizer code operates in error detection mode, can be implemented similarly using this protocol.
We compare the circuit complexity of entanglement distillation between a unitary protocol and our measurement-based protocol. For general stabilizer codes, both circuits scale similarly in the number of two-qubit gates, with an upper bound of $O\left(n\left(n-k\right)\right)$. This is expected, as the unitary decoding circuit derives from the standard form of stabilizer generators [81]. However, stabilizer measurements offer more flexibility. Well-designed codes, such as quantum convolutional codes (Section V) and quantum low-density parity-check (qLDPC) codes (Section VI), support parallel stabilizer measurements, reducing circuit depth. The measurement-based decoder can also decode logical qubits in quantum computing. However, the phase to be corrected is calculated by the sum of single-qubit measurement outcomes, where errors can accumulate. In Section VI, we propose a fault-tolerant protocol for encoding and decoding arbitrary quantum states by combining entanglement distillation, constant-depth decoder, and quantum teleportation.
@@ V Application II: Entanglement Distillation using Quantum Convolutional Codes
Convolutional entanglement distillation implements the stabilizer entanglement distillation protocol using convolutional codes [97, 98, 84, 85, 99, 58]. These codes are widely used in modern communication systems due to their robust error correction capabilities [100] and are key components of Turbo codes [101, 102, 103]. They allow flexible adjustments in code rates to balance error correction strength and efficiency, making them ideal for addressing fluctuating noise. Quantum convolutional codes inherit these benefits and show potential for entanglement distillation. However, they are prone to catastrophic error propagation [86, 87], where a single error can spread indefinitely. To avoid this, encoding circuits are selectively constructed using a restricted set of quantum gates, often requiring many gates and deeper circuits. This increased demand for resources can challenge current quantum technologies. Our measurement-based approach provides an efficient non-catastrophic scheme, making convolutional entanglement distillation viable.
@@
Like block codes, operations on the check matrix correspond to generator operations, such as swapping rows (relabeling generators), adding two rows (multiplying generators), and swapping columns in both submatrices (relabeling qubits). Additionally, multiplying row $g_{i}\left(D\right)$ by $D^{k}$ delays the generator by $k$ frames. Both $g_{i}\left(D\right)$ and $D^{k}g_{i}\left(D\right)$ independently act as generators of the convolutional code. The inverse operator $D^{-k}$ moves the generator forward by $k$ frames. We limit the forward movement to be finite but allow unrestricted delays. Combining shift and addition operations produces convoluted generators, represented by $f\left(D\right)g_{i}\left(D\right)$, where $f\left(D\right)=\sum_{i=d}^{\infty}{f_{i}D^{i}}$ is a Laurent series, $f_{i}$ is a binary number, and $d$ is a lower bounded integer. Note that $f\left(D\right)$ can be expressed as $\frac{p\left(D\right)}{q\left(D\right)}$, where $p\left(D\right)$ and $q\left(D\right)$ are polynomials. For example, $\frac{1}{1+D}=1+D+D^{2}+\cdots$ is a Laurent series. Similarly, $f\left(D^{-1}\right)$ also forms a Laurent series. Two generators $g_{i}\left(D\right)=\left(\bm{x}_{i}\left(D\right)\middle|\bm{z}_{i}\left(D\right)\right)$ and $g_{j}\left(D\right)=\left(\bm{x}_{j}\left(D\right)\middle|\bm{z}_{j}\left(D\right)\right)$ commute if and only if their symplectic inner product is zero, given by
@@
We can verify the commutativity of these operators with the stabilizers using Equation 12. For convenience, we denote $\bm{A}^{\prime}\left(D\right)=\bm{A}_{2}^{T}\left(D^{-1}\right)$, $\bm{C}^{\prime}\left(D\right)=\bm{C}^{T}\left(D^{-1}\right)$, and $\bm{E}^{\prime}\left(D\right)=\bm{E}^{T}\left(D^{-1}\right)$. Generally, the entries of these submatrices are Laurent series, but we assume them to be finite Laurent polynomials. This assumption holds if Gaussian elimination is performed optimally. For cases involving entries as Laurent series, pre-shared Bell states can facilitate decoding with asymptotically zero overhead. These Bell states can be obtained from previous entanglement distillation. For more details, see Appendix C.
@@
Gaussian elimination (see Appendix D for detailed steps) yields the standard form
@@
We compare the circuit complexity of our method with the unitary method described from  [86] for the rate-$\frac{1}{3}$ convolutional code. Our circuit on one side has a depth of $4$, using $12$ two-qubit gates per frame, with the longest two-qubit gate spanning $2$ frames. The unitary decoding circuit, constrained by the non-catastrophic requirement, has a depth of $11$ and $14$ two-qubit gates per frame, with the longest gate spanning $3$ frames. Thus, our method is simpler. While this code is a relatively basic convolutional code, longer convolutional codes can be preferred for better performance, where our method may offer further advantages. However, comparing general convolutional codes remains an open question, requiring theoretical complexity analysis of their unitary encoding circuits.
@@ VI Application III: Efficient Fault-Tolerant Encoding and Decoding for qLDPC Codes
We propose an efficient fault-tolerant protocol for encoding and decoding quantum states in quantum computing, using stabilizer entanglement distillation and quantum teleportation. This scheme is particularly effective for quantum low-density parity-check (qLDPC) codes [78] and surface codes  [79, 80]. The stabilizer generators of these codes can be measured using a constant-depth circuit, as each check operator acts on a constant number of qubits [79, 105, 106, 107]. Logical operators, including single-qubit Hadamard, phase, and CNOT gates, can also be implemented with constant-depth circuits [108, 109, 110]. These combined operations support encoding and decoding circuits with constant depth. Our method surpasses the well-known limitation in encoding surface codes, which states that the depth of the unitary encoding circuit is low bounded by $O\left(log\left(n\right)\right)$ without classical communication [82, 83], where $n$ is the number of physical qubits.
@@
Our protocol involves generating a resource physical-logical Bell state, represented by
@@
All steps operate within a constant circuit depth, enabling efficient encoding and decoding. For example, Figure 5 illustrates stabilizers, logical operators, and decoding measurements for a planar code with parameters $\left[\!\left[2L^{2}-2L+1,1,L\right]\!\right]$. Note that decoding measurements are not unique. Single-qubit measurements can target any logical $X_{L}$ and $Z_{L}$ operators, except for the intersecting qubit, which is identified as the decoded qubit [111].
@@
where $p$ is the error probability of a single-qubit measurement, and $L$ is the lattice length (see Appendix E for a detailed derivation). We can suppress the error by distilling the resource state using a constant-depth circuit. To achieve fault tolerance, the resource state needs only a fidelity of $1-p_{th}$, where $p_{th}$ is the error threshold for fault-tolerant quantum computing with the surface code. This fidelity is attainable with a constant number of distillations, maintaining constant circuit depth. However, successful distillation necessitates $P_{L}<0.5$, defining a threshold for measurement error. Figure 6 shows this threshold as a function of lattice size, following the power law $p\propto L^{-1}$ at large $L$. For an $L=23$ lattice, encoding a logical qubit with $1013$ physical qubits, the measurement error threshold is $0.0196$, which is feasible with current technology.
We discuss the advancements of our results with previous literature. Our encoding and decoding scheme operates on quantum platforms that support multi-dimensional qubit interactions, recently enabled by reconfigurable atom arrays [88, 89]. Łodyga et al. [112] proposed an encoding and decoding scheme for surface codes using stabilizer and single-qubit measurements, asserting that logical operators must intersect at a single physical qubit or at adjacent qubits. However, our findings indicate this condition is not necessary, as our scheme naturally arises from the standard form of stabilizer codes. Unlike their approach, which relies on a specific 3D lattice, our scheme uses entanglement distillation for fault tolerance and is adaptable to various geometries, enhancing its applicability. Bravyi et al. [113] proposed a scheme for preparing logical state $0_{L}$ and logical Bell state of surface codes using measurements within a constant circuit depth. Their scheme requires a prepared $3D$ Raussendorf cluster state [114] as well.

### v2

@@ IV Decoding the Stabilizer Code by Single-Qubit Measurements
We propose a decoder that transforms logical Bell states into physical Bell states using single-qubit measurements and conditional Pauli operators. Since stabilizer and single-qubit measurements can be performed in parallel, this significantly reduces the circuit depth for implementing stabilizer entanglement distillation protocols. We first outline how logical Bell states result from stabilizer measurements on physical Bell states, then explain decoding these logical states into physical states via single-qubit measurements. While presented here for entanglement distillation, this measurement-based decoder applies to any scenario involving decoding stabilizer codes, including quantum computing.
@@
To develop the decoder, we represent the stabilizer code using a binary check matrix $\left(\bm{X}\middle|\bm{Z}\right)$ [91], where $\bm{X}$ and $\bm{Z}$ are submatrices of dimensions $\left(n-k\right)\times n$. The rows represent generators $g_{1}$ through $g_{n-k}$, and the columns of each submatrix correspond to the $n$ physical qubits. In the $i$th row, represented by vector $\left(\bm{x}_{i}\middle|\bm{z}_{i}\right)$, $x_{ij}=1$ indicates a Pauli-$X$ operator on the $j$th qubit, $z_{ij}=1$ indicates a Pauli-$Z$ operator, and $1$ on both indicates a Pauli-$Y$ operator. Two generators $g_{i}$ and $g_{j}$ commute if and only if their symplectic inner product is zero, given by
@@
which follows the same dimensions. The $i$th row, $L_{Z}^{(i)}$, represents the $Z$ operator for the $i$th logical qubit. These logical operators commute with the stabilizer generators. $X_{L}^{(i)}$ and $Z_{L}^{(i)}$ anti-commute, while $X_{L}^{(i)}$ and $Z_{L}^{(j)}$ commute for $i\neq j$.
@@
    The logical operator $X_{L}^{(i)}$ is independent of $X_{j}$ for $j\leq r$ and $n-k<j<n-k+i$, and independent of $Z_{j}$ for $r<j\leq n$. The logical operator $Z_{L}^{(i)}$ is independent of all $X$s and independent of $Z_{j}$ for $r<j<n-k+i$.
@@
$\begin{matrix}g_{1}\otimes g_{1}\;,&Z_{1}\otimes Z_{1}\\ \vdots&\vdots\\ g_{r}\otimes g_{r}\;,&Z_{r}\otimes Z_{r}\\ X_{r+1}\otimes X_{r+1}\;,&g_{r+1}\otimes g_{r+1}\\ \vdots&\vdots\\ X_{n-k}\otimes X_{n-k}\;,&g_{n-k}\otimes g_{n-k}\\ X_{L}^{(1)}\otimes X_{L}^{(1)}\;,&Z_{L}^{(1)}\otimes Z_{L}^{(1)}\\ \vdots&\vdots\\ X_{L}^{(k)}\otimes X_{L}^{(k)}\;,&Z_{L}^{(k)}\otimes Z_{L}^{(k)}\\ \end{matrix}\;.$ (7)
@@
$X_{L}^{(i)}=\left(-1\right)^{\alpha_{x}^{(i)}}X_{n-k+i}\;,$ (10)
where $\alpha_{x}^{(i)}=\sum_{j:C_{ij}^{T}=1}a_{j}+\sum_{j:E_{ij}^{T}=1}a_{r+j}$. Similarly, the logical operator $Z_{L}^{(i)}$ is given by
$Z_{L}^{(i)}=\left(-1\right)^{\alpha_{z}^{(i)}}Z_{n-k+i}\;,$ (11)
where $\alpha_{z}^{(i)}=\sum_{j:\left(A_{2}^{T}\right)_{ij}=1}a_{j}$ and $\left(A_{2}^{T}\right)_{ij}$ is the element of submatrix $\bm{A}_{2}^{T}$ from Equation 5. The single-qubit measurements anti-commute with the code generators, updating the stabilizers in Equation 8 to
$\begin{matrix}\left(-1\right)^{a_{1}}Z_{1}\otimes I\;,&\left(-1\right)^{b_{1}}I\otimes Z_{r}\\ \vdots&\vdots\\ \left(-1\right)^{a_{r}}Z_{r}\otimes I\;,&\left(-1\right)^{b_{r}}I\otimes Z_{r}\\ \left(-1\right)^{a_{r+1}}X_{r+1}\otimes I\;,&\left(-1\right)^{b_{r+1}}I\otimes X_{r+1}\\ \vdots&\vdots\\ \left(-1\right)^{a_{n-k}}X_{n-k}\otimes I\;,&\left(-1\right)^{b_{n-k}}I\otimes X_{n-k}\\ \left(-1\right)^{\alpha_{x}^{(1)}+\beta_{x}^{(1)}}X_{n-k+1}\otimes X_{n-k+1}\;,&\left(-1\right)^{\alpha_{z}^{(1)}+\beta_{z}^{(1)}}Z_{n-k+1}\otimes Z_{n-k+1}\\ \vdots&\vdots\\ \left(-1\right)^{\alpha_{x}^{(k)}+\beta_{x}^{(k)}}X_{n}\otimes X_{n}\;,&\left(-1\right)^{\alpha_{z}^{(k)}+\beta_{z}^{(k)}}Z_{n}\otimes Z_{n}\\ \end{matrix}\;.$
Here, $\bm{b}=b_{1}b_{2}\ldots b_{n-k}$ denotes Bob’s decoding parities from single-qubit measurements, which are independent of Alice’s. $\beta_{x}^{(i)}$ and $\beta_{z}^{(i)}$ are computed in the same way as $\alpha_{x}^{(i)}$ and $\alpha_{z}^{(i)}$, respectively. After the measurements, Alice and Bob discard their first $n-k$ qubits and retain the last $k$ qubits, forming the decoded physical Bell states. They adjust the phases $\alpha_{x}^{(i)}$, $\alpha_{z}^{(i)}$, $\beta_{x}^{(i)}$, and $\beta_{z}^{(i)}$ to 00 by applying Pauli operators to the corresponding qubits, yielding $k$ standard Bell states.
By integrating stabilizer measurements, error correction, and our decoder (see Appendix C for more details), we outline the measurement-based stabilizer entanglement distillation protocol as follows:
@@
This results in $k$ distilled Bell states shared between Alice and Bob. Note that two-way entanglement distillation, where the stabilizer code operates in error detection mode, can be implemented similarly using this protocol. For example, consider the $\left[\!\left[5,1,3\right]\!\right]$ code. Its stabilizer generators are
We can integrate stabilizer measurements and single-qubit measurement decoding into a circuit as shown in Figure 4. This circuit yields both stabilizer parities and decoding parities needed for error correction.
We compare the circuit complexity of entanglement distillation between a unitary protocol and our measurement-based protocol. For general stabilizer codes, both circuits scale similarly in the number of two-qubit gates, with an upper bound of $O\left(n\left(n-k\right)\right)$. This is expected, as the unitary decoding circuit derives from the standard form of stabilizer generators [82]. However, stabilizer measurements offer more flexibility. Well-designed codes, such as quantum convolutional codes (Section V) and quantum low-density parity-check (QLDPC) codes (Section VI), support parallel stabilizer measurements, reducing circuit depth. The measurement-based decoder applies broadly, including in quantum computing, as it relies only on the standard form of logical operators. The logical $X$ and $Z$ operators share commutative Pauli operators on the first $n-k$ qubits, allowing single-qubit measurements on these qubits to preserve logical information. Replacing these Pauli operators with their measurement outcomes decodes the logical operators into single-qubit operators. However, the phase to be corrected is calculated by the sum of single-qubit measurement outcomes, where errors can accumulate. In Section VI, we propose a fault-tolerant protocol for encoding and decoding arbitrary quantum states by combining entanglement distillation, constant-depth decoders, and quantum teleportation.
@@ V Application II: Entanglement Distillation using Quantum Convolutional Codes
Convolutional entanglement distillation implements the stabilizer entanglement distillation protocol using convolutional codes [98, 99, 85, 86, 100, 58]. These codes are widely used in modern communication systems due to their robust error correction capabilities [101] and are key components of Turbo codes [102, 103, 104]. They allow flexible adjustments in code rates to balance error correction strength and efficiency, making them ideal for addressing fluctuating noise. Quantum convolutional codes inherit these benefits and show potential for entanglement distillation. However, they are prone to catastrophic error propagation [102, 87, 88], where a single error can spread indefinitely. To avoid this, encoding circuits are selectively constructed using a restricted set of quantum gates, often requiring many gates and deeper circuits. This increased demand for resources can challenge current quantum technologies. Our measurement-based approach provides an efficient non-catastrophic scheme, making convolutional entanglement distillation viable.
@@
Like block codes, operations on the check matrix correspond to generator operations, such as swapping rows (relabeling generators), adding two rows (multiplying generators), and swapping columns in both submatrices (relabeling qubits). Additionally, multiplying row $g_{i}\left(D\right)$ by $D^{k}$ delays the generator by $k$ frames. Both $g_{i}\left(D\right)$ and $D^{k}g_{i}\left(D\right)$ independently act as generators of the convolutional code. The inverse operator $D^{-k}$ moves the generator forward by $k$ frames. We limit the forward movement to be finite but allow unrestricted delays. Combining shift and addition operations produces convoluted generators, represented by $f\left(D\right)g_{i}\left(D\right)$, where $f\left(D\right)=\sum_{i=d}^{\infty}{f_{i}D^{i}}$ is a Laurent series, $f_{i}$ is a binary number, and $d$ is a lower bounded integer. Note that $f\left(D\right)$ can be expressed as $\frac{p\left(D\right)}{q\left(D\right)}$, where $p\left(D\right)$ and $q\left(D\right)$ are polynomials. For example, $\frac{1}{1+D}=1+D+D^{2}+\cdots$ is a Laurent series. Similarly, $f\left(D^{-1}\right)$ also forms a Laurent series. Two shift-invariant generators $g_{i}(D)=\left(\bm{x}_{i}\left(D\right)\middle|\bm{z}_{i}\left(D\right)\right)$ and $g_{j}(D)=\left(\bm{x}_{j}\left(D\right)\middle|\bm{z}_{j}\left(D\right)\right)$ commute if $g_{i}(D)$ also commutes with the generator $D^{l}g_{j}(D)$. Setting their symplectic inner product to zero for all $l$, the condition for two shift-invariant generators to commute is given by [87]
@@
We can verify the commutativity of these operators with the stabilizers using Equation 12. For convenience, we denote $\bm{A}^{\prime}\left(D\right)=\bm{A}_{2}^{T}\left(D^{-1}\right)$, $\bm{C}^{\prime}\left(D\right)=\bm{C}^{T}\left(D^{-1}\right)$, and $\bm{E}^{\prime}\left(D\right)=\bm{E}^{T}\left(D^{-1}\right)$. Generally, the entries of these submatrices are Laurent series, but we assume them to be finite Laurent polynomials. This assumption holds when Gaussian elimination is performed with row operations using finite Laurent polynomial multipliers (see the example in Appendix E). For cases involving entries as Laurent series, pre-shared Bell states can facilitate decoding with asymptotically zero overhead. These Bell states can be obtained from previous entanglement distillation. For more details, see Appendix D.
@@
Gaussian elimination (see Appendix E for detailed steps) yields the standard form
@@
We compare the circuit complexity of our method with the unitary method described from  [87] for the rate-$\frac{1}{3}$ convolutional code. Our circuit on one side has a depth of $4$, using $12$ two-qubit gates per frame, with the longest two-qubit gate spanning $2$ frames. The unitary decoding circuit, constrained by the non-catastrophic requirement, has a depth of $11$ and $14$ two-qubit gates per frame, with the longest gate spanning $3$ frames. Thus, our method is simpler. While this code is a relatively simple convolutional code, more complex convolutional codes with higher encoding rates and larger code distances may be preferable for improved performance, where our method could offer even greater advantages. However, comparing general convolutional codes remains an open question, requiring theoretical complexity analysis of their unitary encoding circuits.
@@ VI Application III: Efficient Fault-Tolerant Encoding and Decoding for QLDPC Codes
We propose an efficient fault-tolerant protocol for encoding and decoding quantum states in quantum computing, using stabilizer entanglement distillation and quantum teleportation. This approach is particularly effective for quantum low-density parity-check (QLDPC) codes [78], which feature a constant number of qubits per stabilizer generator and a constant number of generators acting on each qubit. This enables parallel stabilizer measurements, significantly reducing the circuit depth compared to unitary encoding methods. In this section, we first present a constant-depth encoder and decoder for surface codes [79, 80, 81], a well-studied class of QLDPC codes, and then demonstrate its fault tolerance. Finally, we discuss the potential and challenges of extending these encoder and decoder protocols to broader QLDPC codes with higher encoding rates.
@@
The efficient encoder and decoder protocol for surface codes involves generating a resource physical-logical Bell state, represented by
@@
Logical operators, such as Pauli, Hadamard, phase, and CNOT gates, can be implemented transversally [106, 107, 108]. Therefore, the entire protocol can operate within constant depth, improving on the well-known lower bound of $\Omega\left(\log\left(n\right)\right)$ for surface code unitary encoding circuits without classical communication [83, 84], where $n$ is the number of physical qubits. For example, Figure 6 illustrates stabilizers, logical operators, and decoding measurements for a planar code with parameters $\left[\!\left[2L^{2}-2L+1,1,L\right]\!\right]$. Note that decoding measurements are not unique. Single-qubit measurements can target any logical $X_{L}$ and $Z_{L}$ operators, except for the intersecting qubit, which is identified as the decoded qubit [109].
@@
where $p$ is the error probability of a single-qubit measurement, and $L$ is the lattice length (see Appendix F for a detailed derivation). We can suppress the error by distilling the resource state using a constant-depth circuit. To achieve fault tolerance, the resource state needs only a fidelity of $1-p_{th}$, where $p_{th}$ is the error threshold for fault-tolerant quantum computing with the surface code. This fidelity is attainable with a constant number of distillations, maintaining constant circuit depth. However, successful distillation necessitates $P_{L}<0.5$, defining a threshold for measurement error. Figure 7 shows this threshold as a function of lattice size, following the power law $p\propto L^{-1}$ at large $L$. For an $L=23$ lattice, encoding a logical qubit with $1013$ physical qubits, the measurement error threshold is $0.0196$, which is feasible with current technology.
Noisy stabilizer measurements present another challenge, potentially introducing errors during correction and leading to logical errors. This issue can be addressed by multiple measurement rounds [110, 111] or by preparing high-quality entangled states [106, 112, 113]. Our protocol uses a constant number of stabilizer measurement rounds, with remaining errors corrected through entanglement distillation of the resource state $\left|\Phi_{R}\right\rangle$. Typically, $\Theta(d)$ rounds of stabilizer measurements are required to suppress logical errors to $p^{\Omega{(d)}}$ [106, 108], where $d$ is the surface code distance and $p$ is the measurement error probability. However, for the encoder, this requirement can be relaxed, as logical errors on Bob’s part of $\left|\Phi_{R}\right\rangle$ can be transferred to Alice’s qubit via the Bell-state identity (see Appendix A). These errors can then be interpreted as errors on the input states, primarily determined by very few physical gates. Consequently, the fidelity of the resource state only needs to meet the fault-tolerance threshold, supporting a constant number of rounds for stabilizer measurements and distillation.
Quantum LDPC codes with constant encoding rates offer advantages over 2D surface codes [114, 78]. Our protocol is compatible with these codes at each step, though some challenges remain. One issue is handling non-local stabilizer measurements, as qubits within the same stabilizer may be spatially separated. According to the BPT bound [115], the best QLDPC codes with local stabilizers in $D$ dimensions are constrained by parameters $\left[\!\left[n,k=n^{1-\frac{2}{D}},d=n^{1-\frac{1}{D}}\right]\!\right]$ [116, 117]. Surface codes are optimal for $D=2$ but achieve only a rate of $1/n$, while asymptotically constant-rate QLDPC codes become feasible in $D\geq 3$ [118, 119, 120]. Multi-layer architectures enable higher-dimensional connections [121, 122], and reconfigurable neutral atoms support long-range connectivity [90]. However, implementing non-local stabilizer measurements with these atoms may require atom rearrangement, potentially increasing gate times. The single-shot property, which enables quantum error correction with a single round of syndrome measurements, is often needed for the efficient implementation of QLDPC codes. Certain QLDPC codes, including specific topological codes [110, 123], quantum expander codes [124, 90], and quantum Tanner codes [125], are currently known to exhibit this property. Another challenge is the limitation of transversal gates, as each block encodes multiple qubits, complicating selective logical operations. Research on logical gates for QLDPC codes is ongoing. For example, code deformation [126] provides a universal gate set through state injection within a single block. Another approach uses state teleportation into ancilla codes, enabling logical operations by separating computation and memory [90, 127, 128, 129, 130]. However, these logical gates may increase circuit depth. The challenges discussed above are general to quantum computing with QLDPC codes rather than specific to our protocol [131]. Our protocol remains more efficient than unitary encoding, benefiting from QLDPC code sparsity and classical communication. For general stabilizer codes, both unitary encoding and our measurement-based encoding have a circuit depth of $O(n)$ and a gate count of $O(n^{2})$ [82].
[…]